Solve the equation (5-2x)^2 = 25 by using the Square Root Property
Answer by edjones(8007)

(5-2x)^2=25
5-2x=+-5 sqrt of both sides.
-2x=-5+-5
x=(-5+-5)/-2
x=0/-2, x=-10/-2
x=0, x=5
Always check:
5^2=25 true
(5-10)^2=25 true
Ed